OpenStudy (anonymous):

An uncharged capacitor and a resistor are connected in series to a source of EMF. If ε = 6.63 V, C = 21.1 μF, and R = 122 Ω, calculate the time constant τ of the circuit.

OpenStudy (anonymous):

Isn't the time constant for resistor capacitor circuits: \[\tau = RC\] Where you need to be consistent with your units. If you have ohms you need to use Farads. If you have microFarads you need to use megaOhms.

OpenStudy (anonymous):

I am still a little confused

OpenStudy (anonymous):

\[1ohm = \frac{ kg \times m^2 }{ s^3 \times A^2 }\] \[1 Farad = \frac{ s^4 \times A^2 }{ m^2 \times kg }\]

OpenStudy (anonymous):

If you multiply them together your units will all disappear except for Seconds.

OpenStudy (anonymous):

You currently have microfarads and ohms. You either need to change convert to just Farads which would be easier or to megaOhms so that you are consistent.

OpenStudy (anonymous):

so now converted to 2.11E-5 and 122.06

OpenStudy (anonymous):

You only need to convert 1

OpenStudy (anonymous):

and I just multiply them together?

OpenStudy (anonymous):

You need to have Ohms times Farads or microFarads times megaOhms.

OpenStudy (anonymous):

You currently have microFarads and Ohms.

OpenStudy (anonymous):

okay so I converted 21.1 to 2.11E-5 F then I times that by the 122ohms?

OpenStudy (anonymous):

Okay I got that to be 2.57ms which is right, now would you know how to find the max charge?

OpenStudy (anonymous):

"find the maximum charge on the capacitor"

OpenStudy (anonymous):

nvm I figured that out lol. But last one

OpenStudy (anonymous):

Calculate the charge on the capacitor after one time constant.

OpenStudy (anonymous):

I don't recall the formulas so I have to look them up but glad its getting easier now. \[Q = CV_b(1-e^{-\frac{ t }{ RC }})\] Where Vb represents the voltage of the battery

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

My pleasure